OpenStudy (anonymous):

Determine the center of the circle given by the equation x^2-8x+y^2-10y+5=0

OpenStudy (anonymous):

Complete the squares in x and y.

OpenStudy (anonymous):

complete the square via \[x^2-8x+y^2-10y=-5\] \[(x-4)^2+(y-5)^2=-5+4^2+5^2=36\] center is (4,5) radius is \[r=\sqrt{36}=6\]